1. Solve the following inequlity 2x/x-2>x+4
1
Expert's answer
2019-06-06T09:58:01-0400

"\frac{2x}{x-2}>x+4"

If "x>2:\;\;2x>(x-2)(x+4)" or "x^2<8" or "2<x<2\sqrt{2}" .

If "x<2:\;\;2x<(x-2)(x+4)" or "x^2>8" or "-\infty<x<-2\sqrt{2}"

Answer: "2<x<2\sqrt{2}\;\;or\;\;-\infty<x<-2\sqrt{2}"